# ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning

### v3

@@ C.1 Prompt Templates for Filtering Explicit Personae
C.2 Prompt Templates for Implicit Feature Ranking
We provide our hand-crafted prompt template for ranking implicit personae opinions in the usefulness order below:
@@ C.2 Prompt Templates for Implicit Feature Ranking
C.3 Prompt Templates for Baselines Techniques
We use the same prompt templates for ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b). The template prompts for baselines are presented below.
C.3.1 W/o Persona Santurkar et al. (2023)
The W/o Persona prompt is provided below.
@@ C.3 Prompt Templates for Baselines Techniques
C.3.3 Self-refine Madaan et al. (2023)
The Self-refine prompts Madaan et al. (2023) are provided below, feedback step and refine step respectively.
@@
You are given a question, an answer to that question and a feedback to the answer. Based on the feedback, refine your answer and generate the final answer in around 170 words. Question: {test_question} Answer: {selected_choice} Feedback: {feedback} Refined answer: new_choice + explanation
C.3.4 Chain-of-Thought Kojima et al. (2022)
The CoT prompt template is provided below.
@@
C.3.5 Chain-of-Opinion (Ours)
Our CoO prompt template is provided below.
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Based on the above information, answer the following question step-by-step by explaining and analyzing each of the Opinions and Demographic Information: Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A. or B. or C. or D. or E.... Explanations:...
C.4 Prompting Costs for API Models
Our prompting costs for API models are reported in table 10. We observe that for ChatGPT and ChatGPT-Instruct, ChOiRe costs around $7$ and $10$ more US\$ dollars in total compared to baseline DIO-top8 + SC. However, these extra amounts of costs are worth it because we gain significant improvements over all the baselines and especially huge improvements for some topics. Additionally, for GPT-4, ChOiRe costs a similar price with the baseline DIO-top8 while DIO-top8 + SC costs nearly double our price. This is because we perform the FEA and LLMtop-$K$ steps of ChOiRe by ChatGPT, which are relatively cheap.
DIO-top8 DIO-top8 + CoT DIO-top8 + SC ChOiRe Model
Ave. consumed #tokens 562.72 623.62 995.89 3142.86 ChatGPT
Total US\$ 3.01 3.73 6.82 13.95 ChatGPT
Ave. consumed #tokens 562.72 630.58 1019.31 3121.72 ChatGPT-Instruct
Total US\$ 3.12 3.84 7.11 19.99 ChatGPT-Instruct
Ave. consumed #tokens 559.27 - 1021.14* 3180.82 GPT-4
Total US\$ 91.19 - 226.15* 123.30 GPT-4
Table 10: Prompting cost analysis of ChOiRe and other baselines as of 1st Feb 2024. * denotes our estimation on $50$ random samples.
@@ Appendix D Human Evaluation
D.1 Human Rating Instructions
Our details of human rating instructions are provided in table 11 for all the criteria. It is worth noting that selecting all features can’t get a high FEA Satisfaction score, according to our instructions. In addition, if the selected explicit personae fall among several scores, the annotators are instructed to take the minimum score.
@@ D.1 Human Rating Instructions
1: The CoO has limited or flawed reasoning thoughts with inadequate support.
CoO Reasonableness 2: The CoO has some reasoning thoughts with decent support but room for improvement.
3: The CoO has strong, clear, and well-supported reasoning thoughts with a comprehensive understanding.
1: The generated CoO explanation does not mention more than $4$ attributes/opinions from explicit and implicit personae.
CoO Follow the Instruction 2: The generated CoO explanation somewhat follows the instruction by involving more than $4$ attributes/opinions but room for improvement.
3: The generated CoO explanation follows perfectly the instruction via explaining all the explicit and implicit attributes one by one.
@@
Appendix E Examples
E.1 FEA Example with ChatGPT
@@
E.2 Example of High Disagreement between Rankings
@@
E.3 Example of Inconsistent Answers Generated by CoT
@@
E.4 Example of Chain of Opinion Reasoning
Fig. 11 presents an example of the answer generated by ChatGPT using Chain of Opinion (ours) versus Chain of Thought Wei et al. (2022) prompting methods.
Figure 11: Example of an answer generated by Chain of Opinion versus Chain of Thought prompting with ChatGPT.
E.5 Example of Answer Consistency with Dynamic Numbers of Opinions
Fig. 12 shows an example of the answer generated by GPT-4 using Chain of Opinion (ours) reasoning with different numbers of provided historical opinions.
Figure 12: Example of our answer consistency technique (ours), generated by GPT-4.
E.6 Error Analysis Examples
Fig. 13 illustrates our error analysis examples of ChOiRe with ChatGPT. The top-left frame is an example of FEA missing key explicit personae. The bottom one is an instance demonstrating the error of the LLMtop-$K$ algorithm including less relevant opinions. The top-right rectangular is an example from GPT-4, showing that it does not follow human instructions to predict opinion via chain-of-opinion reasoning.
Figure 13: Error analysis examples of ChOiRe with ChatGPT.

### v5

@@ C.2 Prompt Template for COO’s FEA
C.3 Prompt Template for COO’s LLMtop-K
@@
C.4 Prompt Template for COO’s VBN Reasoning
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Given the following question: Question: {question} Answer choices: {choice} Answer the above question by following the steps below: Analyze the user’s demographics and ideology one by one to infer their social and environmental values. Wrap this analysis by <EV> and </EV>. Analyze the user’s historical opinions to infer their beliefs and norms from their social and environmental values. Wrap this analysis by <PBN> and </PBN>. From the above analyses, which opinion he is likely to choose? Answer: A. or B. or C. or D. or E....
C.5 Prompt Templates for Baselines
We use the same prompt templates for ChatGPT OpenAI (2022), ChatGPT-it OpenAI (2023a), GPT-4 OpenAI (2023b). The template prompts for baselines are presented below.
C.5.1 W/o Persona Santurkar et al. (2023)
@@ C.5 Prompt Templates for Baselines
C.5.3 Self-refine Madaan et al. (2023)
@@
You are given a question, an answer to that question, and feedback to the answer. Based on the feedback, refine your answer and generate the final answer in around 170 words. Question: {test_question} Answer: {selected_choice} Feedback: {feedback} Refined answer: new_choice + explanation
C.5.4 Chain-of-Thought Kojima et al. (2022)
@@ Appendix D Human Evaluation
D.1 Human Grading Instructions
Our details of human rating instructions are provided in Tab. 12 for all the criteria. It is worth noting that selecting all features can’t get a high FEA Satisfaction score, according to our instructions. In addition, if the selected explicit personae fall among several scores, the annotators are instructed to take the minimum score.
@@ D.1 Human Grading Instructions
1: The VBN has limited or flawed values, beliefs, norms thoughts with inadequate support.
VBN Reasonableness 2: The VBN has some values, beliefs, norms thoughts with decent support but room for improvement.
3: The VBN has strong, clear, and well-supported values, beliefs, norms thoughts with a comprehensive understanding.
1: The generated VBN explanation does not mention more than $4$ attributes/opinions from explicit and implicit personae.
VBN FI 2: The generated VBN explanation somewhat follows the instruction by involving more than $4$ attributes/opinions but room for improvement.
3: The generated VBN explanation follows perfectly the instructions via explaining all the explicit and implicit attributes one by one.
@@
Appendix E Error Analyses and Examples
E.1 Error Analyses
FEA misses key explicit personae.
Despite showing promising results with FEA, LLMs sometimes misselect relevant personae. One such example is the top-left of Section E.2. We observe that in this case, our annotators can’t grade a high FEA satisfaction score because "Education" and "Age" are also two important personae as they can influence one’s understanding of workplace dynamics significantly, which are deselected by ChatGPT.
LLMtop-$K$ opinions consist of less relevant ones.
We observe LLMs frequently include less relevant, or even irrelevant opinions to the ranked list such as in Section E.2-bottom. We attribute this to the challenge of this task, even for humans it might require substantial cognitive effort.
LLMs may not follow the instructions to perform VBN reasoning.
Although ChatGPT and ChatGPT-it demonstrate a robust ability to perform VBN reasoning (Tab. 3), the same level of proficiency is not observed in Mistral and GPT-4, as exampled in Section E.2-top-right. We posit this disparity arises from the fact that ChatGPT and ChatGPT-it excel in comprehending and executing human instructions, while GPT-4 excels primarily in generating coherent text.
E.2 Examples
FEA example with ChatGPT.
@@ E.2 Examples
Example of high disagreement between rankings.
@@
Example of inconsistent answers generated by CoT.
@@
Human evaluation example.
Fig. 8 illustrates our human evaluation example of COO with ChatGPT. The top-left frame is an example of FEA missing key explicit personae. The bottom one is an instance demonstrating the error of the LLMtop-$K$ algorithm including less relevant opinions. The top-right rectangular is an example from GPT-4, showing that it does not follow human instructions to predict opinion via chain-of-opinion reasoning.
Figure 8: Error analysis examples of COO with ChatGPT.